OpenStudy (anonymous):

A mechanic pushes a(n) 2960 kg car from rest to a speed of v, doing 5023 J of work in the process. During this time, the car moves 20.9 m. Neglect friction between car and road. Find the speed v Then, find the horizontal force exerted on the car.

OpenStudy (anonymous):

5023 = 0.5(2960)v^2 v = [(5023)/(0.5)(2960)] ^(0.5)

OpenStudy (anonymous):

The horizontal force = 5023 / 20.9